OpenStudy (anonymous):

solve: m/m-7=(m/m-7) + 11

OpenStudy (anonymous):

is it (m-7) or m/m - 7

OpenStudy (anonymous):

because m/m = 1 so i guess it's m/(m-7)

OpenStudy (anonymous):

m/m-7

OpenStudy (anonymous):

Sure of how it's written? Add more parenthesis if needed... 1-6-11=(m/m-7) -16(m-7)=m -15m=-112 m=112/15

OpenStudy (anonymous):

positive

OpenStudy (anonymous):

no real solutions?

OpenStudy (anonymous):

m/ (m-7)

OpenStudy (anonymous):

hold on give me a sec

OpenStudy (anonymous):

im getting a gross looking polynomial and i don't have my calc :'(

OpenStudy (anonymous):

oh wait

OpenStudy (anonymous):

can you just subtract the (m/m-7)?

OpenStudy (anonymous):

you could and that's why i think it might me no solution

OpenStudy (anonymous):

NOPE :P

OpenStudy (anonymous):

did you try multiplying by (m-7)?

OpenStudy (anonymous):

for it to cancel wouldn't you need to multiply by (m+7)

OpenStudy (anonymous):

both sides yes

OpenStudy (anonymous):

then you et -7 =m

OpenStudy (anonymous):

*get

OpenStudy (anonymous):

YAY :D

OpenStudy (anonymous):

i'll do it out! awesome thanks!

OpenStudy (anonymous):

NOOOO!